### v1

# Dual-Stage Cross-Modal Network with Dynamic Feature Fusion for Emotional Mimicry Intensity Estimation

Abstract
Emotional Mimicry Intensity (EMI) estimation serves as a critical technology for understanding human social behavior and enhancing human-computer interaction experiences, where the core challenge lies in dynamic correlation modeling and robust fusion of multimodal temporal signals. To address the limitations of existing methods in insufficient exploitation of modal synergistic effects, noise sensitivity, and limited fine-grained alignment capabilities, this paper proposes a dual-stage cross-modal alignment framework. First, we construct vision-text and audio-text contrastive learning networks based on an improved CLIP architecture, achieving preliminary alignment in the feature space through modality-decoupled pre-training. Subsequently, we design a temporal-aware dynamic fusion module that combines Temporal Convolutional Networks (TCN) and gated bidirectional LSTM to respectively capture the macro-evolution patterns of facial expressions and local dynamics of acoustic features. Innovatively, we introduce a quality-guided modality fusion strategy that enables modality compensation under occlusion and noisy scenarios through differentiable weight allocation. Experimental results on the Hume-Vidmimic2 dataset demonstrate that our method achieves an average Pearson correlation coefficient of 0.35 across six emotion dimensions, outperforming the best baseline by 40%. Ablation studies further validate the effectiveness of the dual-stage training strategy and dynamic fusion mechanism, providing a novel technical pathway for fine-grained emotion analysis in open environments.
**footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
1 Introduction
Emotional mimicry stands as one of the core behaviors in human social interactions, where individuals unconsciously replicate others’ facial expressions, vocal tones, or body movements to rapidly establish emotional resonance and strengthen social bonds. This psychological mechanism not only holds significant implications for understanding human social behavior but also provides natural inspiration for affective computing technologies in artificial intelligence. With the rapid development of applications such as virtual assistants, emotional robots, and mental health assessment, enabling machines to accurately perceive and quantify dynamic changes in human emotions has become a central challenge in human-computer interaction research. Among these, automatic estimation of Emotional Mimicry Intensity (EMI), capable of capturing the continuity and nuanced variations in emotional transmission, has gradually emerged as a cutting-edge direction in affective computing.
However, existing methods still face significant bottlenecks in addressing EMI estimation tasks. First, emotional mimicry is inherently a multimodal temporal process – subtle changes in visual expressions and dynamic fluctuations in vocal prosody often intertwine nonlinearly. Current research predominantly employs unimodal analysis (e.g., facial video only) or simplistic feature concatenation strategies, resulting in underutilized inter-modal synergy. For instance, baseline methods using ViT+GRU for visual information achieve merely 0.09 average Pearson correlation coefficient, indicating that unimodal visual models struggle to capture temporal evolution patterns of emotional intensity. Second, fine-grained alignment of emotional intensity requires cross-modal semantic understanding capabilities. Although Wav2Vec2 audio features combined with linear layers attain a relatively higher performance of 0.24, their unimodal nature limits robustness in complex scenarios (e.g., strong noise, occlusion). Furthermore, existing fusion strategies (e.g., simple average fusion in baselines with only 0.25 improvement) fail to account for differential information contributions across modalities at varying emotional phases, leaving multimodal potential underdeveloped.
To address these challenges, this study proposes a multimodal fusion-based EMI estimation framework with three core contributions: First, we implement a dual-stage training paradigm. During the first stage, we construct image-text and audio-text dual-path alignment networks based on the CLIP architecture[18], leveraging its contrastive learning mechanism to achieve primary alignment between text-visual and text-audio modalities. In the second stage, a modality mapping layer facilitates secondary alignment across image-text-audio triples. Compared to conventional approaches that separately extract features using three independent modality encoders, CLIP’s alignment properties enable feature extraction that preserves cross-modal characteristics. Second, we design quality-aware dynamic fusion to adaptively reduce the contribution weights of impaired modalities. Third, temporal features are extracted through a Temporal Convolutional Network (TCN)[1] with segment-wise average pooling to suppress noisy frame interference and generate robust global representations, while audio modality employs bidirectional LSTM[19] to encode spectral-temporal dependencies.
This study is based on the EMI track of the 8th ABAW (Affective Behavior Analysis in-the-wild) Competition[11, 10, 5, 8, 17, 9, 14, 15, 6, 7, 12, 2, 3, 13, 4, 16]. The adopted dataset, Hume-Vidmimic2, comprises 15,000 videos totaling over 25 hours of content.Participants in these videos mimic emotional expressions observed in seed videos, with their performance subsequently evaluated across multiple dimensions including ”admiration,” ”amusement,” ”determination,” ”empathic pain,” ”excitement,” and ”joy.”
Through systematic experiments on the Hume-Vidmimic2 dataset, this paper validates the effectiveness of the proposed method. Compared to baseline models, our framework achieves a breakthrough performance with an average Pearson correlation coefficient of 0.35 across six emotional dimensions, demonstrating 40% improvement over the best baseline (0.25).
The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 details the employed multimodal features and model architecture. Section 4 describes experimental implementation specifics and presents result analyses. Finally, Section 5 concludes our work.
2 Related Work
2.1 Multimodal Emotion Recognition Techniques
Multimodal emotion recognition enhances the robustness of affective analysis by fusing complementary information from visual, audio, and textual modalities. Early work such as Dynamic Fusion Network (DFN) and Tensor Fusion Network (TFN)[20, 21], which dynamically weights features through inter-modal gating mechanisms, yet its static weight allocation struggles to adapt to temporal variations in emotional intensity. Recent studies focus on cross-modal interaction modeling: Tsai et al. (2019) [9] designed a Memory-Augmented Cross-modal Attention (MULT) that utilizes external memory banks to store modality-shared semantics but fails to explicitly address semantic discrepancies caused by modal heterogeneity. Hazarika et al. (2020) [10] introduced cross-modal contrastive loss to narrow inter-modal distribution gaps via adversarial training, though their single-stage alignment strategy cannot handle hierarchical correlations in multimodal data. Compared to these approaches, our method decouples cross-modal semantic alignment through a dual-stage contrastive alignment mechanism, effectively mitigating inter-modal information redundancy and conflicts.
2.2 Temporal Modeling and Positional Encoding
The temporal dynamics of emotional mimicry intensity demand models capable of capturing both long-range dependencies and local contextual patterns. For visual temporal modeling, Lea et al. (2017) [11] proposed Temporal Convolutional Networks (TCN), which capture long-range dependencies between video frames through stacked dilated convolutions. However, their fixed receptive fields struggle to adapt to the non-stationary characteristics of emotional intensity variations. In audio temporal modeling, Hershey et al. (2021) [12] extracted spectral features via Wav2Vec2.0’s self-supervised pretraining but neglected explicit modeling of phoneme-emotion intensity mapping relationships. Regarding positional encoding, Vaswani et al. (2017) [13] introduced absolute positional encoding in Transformers, yet its adaptability for multimodal temporal alignment remains limited. Huang et al. (2020) [14] developed Modality-Adaptive Relative Positional Encoding (MARPE), which distinguishes video, audio, and textual characteristics through learnable parameters to facilitate differential modality fusion. Nevertheless, this approach overlooks the impact of modality quality fluctuations on fusion effectiveness. To address these limitations, we design a modality-aware dynamic fusion mechanism that automatically reduces the contribution weights of impaired modalities when quality discrepancies exist in multimodal signals (e.g., facial occlusion or environmental noise). Our framework combines TCN and bidirectional LSTM to respectively model local-global temporal dependencies in visual and audio modalities, while employing segment-wise average pooling to suppress interference from noisy frames.
2.3 Cross-Modal Alignment Techniques
Cross-modal alignment aims to establish shared semantic spaces for multimodal data. Radford et al. (2021) [1] proposed the CLIP model, which achieves image-text semantic alignment through contrastive learning, yet without extending to the audio modality. Wu et al. (2022) [15] constructed an audio-text joint embedding space (CLAP) based on the CLIP architecture, but their single-stage alignment strategy fails to balance multimodal semantic consistency. Recent work like Alayrac et al. (2022) [16] introduced the Flamingo model, achieving fine-grained image-text alignment through cross-attention, though its computational complexity limits practical deployment in real-time tasks. In affective computing, Yang et al. (2023) [17] aligned visual-audio-text features using contrastive learning but failed to explicitly model cross-modal uncertainty. This paper proposes a dual-stage alignment framework: The first stage independently aligns image-text and audio-text pairs through CLIP and CLAP architectures, while the second stage introduces a modality mapping layer to optimize triple-modality joint semantics, significantly enhancing fine-grained consistency in emotional intensity estimation.
3 Method
The proposed Dynamic Multimodal Emotional Mimicry Intensity (EMI) Estimation Framework adopts a dual-stage training paradigm, integrating cross-modal alignment with hierarchical fusion mechanisms to address challenges in multimodal temporal modeling, dynamic weight allocation, and fine-grained semantic alignment. A detailed exposition follows from four technical perspectives: feature extraction, alignment strategies, fusion design, and training optimization.
3.1 Multimodal Data Preprocessing
The input data comprises three modalities: visual, audio, and textual. The preprocessing methods for each modality are detailed below.
[…]
Audio Modality: Raw audio signals are segmented into 20ms windows strictly aligned with video frames. Acoustic features are extracted using the pretrained Wav2Vec2.0 model, producing 1024-dimensional feature sequences $A=\{a_{1},a_{2},...,a_{T}\}$. The model’s self-supervised pretraining on unlabeled speech data enhances its representational capacity for emotion-related acoustic attributes like pitch and speaking rate.
Text Modality: The text processing pipeline begins with speech-to-text conversion using the Whisper model , which employs an end-to-end deep learning architecture optimized for speech recognition. Unlike conventional ASR systems, Whisper innovatively adopts a multi-task training paradigm, simultaneously optimizing speech recognition, language identification, and voice activity detection across 1.28 million hours of diverse audio data. This approach significantly improves robustness in noisy environments and multilingual adaptability. Specifically, we utilize a pruned and fine-tuned version, Whisper-large-v3-turbo , for high-accuracy transcription to obtain semantically complete utterances $s$.
Transcribed texts are encoded through the pretrained BERT model  using its 12-layer Transformer architecture to extract 768-dimensional context-aware feature vectors $f_{s}$. As a foundational NLP model, BERT’s dual pretraining objectives (masked language modeling and next-sentence prediction) enable effective capture of deep semantic relationships, providing reliable linguistic representations for our framework.
3.2 Overall Framework
The proposed multimodal EMI estimation framework incorporates a dual-stage training paradigm and multimodal fusion architecture. The first stage achieves independent semantic alignment between image-text and audio-text pairs through contrastive learning. The second stage constructs cross-modal temporal modeling and dynamic fusion modules while freezing modality encoders. Final emotional mimicry intensity estimation is implemented end-to-end via a multi-layer Transformer encoder.
3.3 Dual-Stage Cross-Modal Alignment
Stage I: Modality-Text Contrastive Alignment Building upon the CLIP architecture [radford2021learning], we establish dual-path contrastive learning networks for vision-text and audio-text pairs. For visual modality, given video frame sequence $V=\{v_{t}\}_{t=1}^{T}$, frame-wise features $f_{v}^{t}\in\mathbb{R}^{d_{v}}$ are extracted through a ViT encoder, with segment-wise average pooling yielding video-level global features $\bar{f}_{v}$. Simultaneously, textual modality employs a pretrained BERT encoder to map transcribed speech $S$ into text features $f_{s}\in\mathbb{R}^{d_{s}}$. The vision-text contrastive loss is defined as:
$\mathcal{L}_{v2t}=-\log\frac{\exp(\bar{f}_{v}\cdot f_{s}/\tau)}{\sum_{i=1}^{N}\exp(\bar{f}_{v}\cdot f_{s_{i}}/\tau)}$ (1)
where $\tau$ denotes the temperature coefficient and $N$ the batch size. For audio modality, the CLAP architecture  processes acoustic signals $A$ through Wav2Vec2.0  to extract spectral features $f_{a}\in\mathbb{R}^{d_{a}}$, with analogous audio-text contrastive loss $\mathcal{L}_{a2t}$ following Equation (1).
Stage II: Tri-Modal Joint Optimization With frozen first-stage vision/audio/text encoders, we construct cross-modal mapping layers. Video features $\{f_{v}^{t}\}$ are processed by Temporal Convolutional Networks (TCN) to capture multi-scale temporal dependencies through stacked dilated convolutions:
$h_{v}^{t}=\text{TCN}(f_{v}^{t};\{\mathbf{W}_{k}^{d_{k}}\}_{k=1}^{K})$ (2)
where $d_{k}$ indicates the dilation rate of the $k$-th convolutional layer and $K$ the total layer count. Audio features $\{f_{a}^{t}\}$ are fed into bidirectional LSTM to model spectral-temporal dynamics:
$h_{a}^{t}=\text{BiLSTM}(f_{a}^{t};\mathbf{W}_{f},\mathbf{W}_{b})$ (3)
Through modality mapping layers (MLP), $h_{v}^{t}$, $h_{a}^{t}$ and text features $f_{s}$ are projected into a shared space and concatenated along temporal dimension:
$H_{\text{fuse}}^{t}=[\text{MLP}(h_{v}^{t});\text{MLP}(h_{a}^{t});\text{MLP}(f_{s})]$ (4)
3.4 Temporal Feature Enhancement
Visual Modality Processing: Building upon TCN outputs $h_{v}^{t}$, we employ a segment-wise average pooling strategy: partition videos into $M$ equal-length segments $\{C_{m}\}_{m=1}^{M}$, then compute per-segment feature means:
$\hat{h}_{v}^{m}=\frac{1}{|C_{m}|}\sum_{t\in C_{m}}h_{v}^{t}$ (5)
This operation suppresses local noise while preserving critical frame information for emotional intensity evolution.
Audio Modality Processing: A gated attention mechanism is applied to BiLSTM outputs $h_{a}^{t}$:
$\alpha_{a}^{t}=\sigma(\mathbf{W}_{g}[h_{a}^{t};\Delta h_{a}^{t}])$ (6)
where $\Delta h_{a}^{t}=h_{a}^{t}-h_{a}^{t-1}$ denotes adjacent-frame feature differences, and $\sigma$ represents the sigmoid function. The enhanced features are calculated as $\tilde{h}_{a}^{t}=\alpha_{a}^{t}\odot h_{a}^{t}$.
3.5 Quality-Aware Dynamic Fusion
To mitigate the impact of inevitable noise in individual modalities, we design a modality quality assessment module for dynamic adjustment of multimodal contribution weights. For timestep $t$, modality quality scores are computed as:
$\ q_{m}=\text{MLP}(h_{m}),\ m\in\{v,a,s\}\$ (7)
Dynamic fusion weights are generated via softmax normalization:
$[\beta_{v}^{t},\beta_{a}^{t},\beta_{s}^{t}]=\text{softmax}([q_{v}^{t},q_{a}^{t},q_{s}])$ (8)
The weighted fused features are fed into a multi-layer Transformer encoder for emotional intensity prediction:
$\hat{y}^{t}=\text{Transformer}(\beta_{v}^{t}\hat{h}_{v}^{m}+\beta_{a}^{t}\tilde{h}_{a}^{t}+\beta_{s}^{t}f_{s})$ (9)
This design enables adaptive attenuation of impaired modalities’ contributions when quality discrepancies exist in multimodal signals (e.g., facial occlusion or environmental noise).
3.6 Training Strategy
We adopt a two-stage training paradigm: Stage I optimizes modality-text alignment using contrastive loss $\mathcal{L}_{v2t}+\mathcal{L}_{a2t}$; Stage II freezes encoder parameters and optimizes regression tasks through mean squared error loss $\mathcal{L}_{\text{MSE}}=\frac{1}{T}\sum_{t=1}^{T}(y^{t}-\hat{y}^{t})^{2}$.
4 Experiment
4.1 Datasets and Experimental Setup
Experiments are conducted on the Hume-Vidmimic2 multimodal affective dataset, which contains 15,000 video samples spanning six emotional dimensions (”Admiration”, ”Amusement”, ”Determination”, ”Empathic Pain”, ”Excitement”, ”Joy”). Each sample’s emotional intensity score, ranging within $[0,1]$, is annotated by professional annotation teams following standardized protocols. The dataset is partitioned into training, validation, and test sets. Statistical details of each split are provided in Table 1.
Partition Duration Samples
Train 15:07:03 8072
Validation 9:12:02 4588
Test 9:04:05 4586
$\sum$ 33:23:10 17246
Table 1: Hume-Vidmimic2 partition statistics.
4.2 Implementation Details
Evaluation Metrics The mean Pearson correlation coefficient ($\rho$) serves as the primary evaluation metric for intensity estimation, measuring the linear correlation between predicted emotional response intensities and ground truth values. This metric is defined as:
$\rho=\sum_{i=1}^{6}\frac{\rho_{i}}{6}$
where $\rho_{i}\ (i\in\{1,2,\cdots,6\})$ denotes the correlation coefficient for each of the six emotion categories, calculated as:
$\rho_{i}=\frac{\text{cov}(y_{i},\hat{y_{i}})}{\sqrt{\text{var}(y_{i})\text{var}(\hat{y_{i}})}}$
Here: $\text{cov}(y_{i},\hat{y_{i}})$ represents the covariance between predicted and target values.$\text{var}(y_{i})$ and $\text{var}(\hat{y_{i}})$ denote the variances of target and predicted values respectively.
4.3 Training Strategy Optimization
Cosine Annealing Scheduling: During the contrastive pretraining phase (50 epochs), we implement Stochastic Gradient Descent with Warm Restarts (SGDR) with learning rate updates governed by:
$\eta_{t}=\eta_{\min}+\frac{1}{2}(\eta_{\max}-\eta_{\min})\left(1+\cos\left(\frac{T_{\text{cur}}}{T_{i}}\pi\right)\right)$ (9)
where $\eta_{\max}=1\times 10^{-5}$, $\eta_{\min}=1\times 10^{-7}$, $T_{i}$ denotes cycle length (10 epochs), and $T_{\text{cur}}$ counts iterations within current cycle. Each cycle concludes with learning rate resetting while preserving optimizer momentum via warm restarts to accelerate convergence.
EMA Exponential Smoothing: Parameter-level Exponential Moving Average (EMA) is activated during fine-tuning with momentum decay factor $\gamma=0.999$. The update rule follows:
$\theta_{\text{EMA}}^{(t)}=\gamma\theta_{\text{EMA}}^{(t-1)}+(1-\gamma)\theta^{(t)}$ (10)
where $\theta^{(t)}$ represents model parameters at step $t$, and $\theta_{\text{EMA}}^{(t)}$ denotes smoothed parameters. Validation-phase inference employs EMA parameters to suppress training oscillations and enhance test-set generalization.
5 Results
5.1 Unimodal Performance Analysis
As shown in Table 3, our enhanced unimodal models demonstrate consistent improvements over baseline implementations across both modalities. For the visual modality, it can be observed that our ViT model pre-trained using contrastive learning demonstrates significantly superior performance compared to the official ViT features. our modified ViT architecture achieves a Pearson correlation coefficient ($\rho_{val}$) of 0.16, representing a 77.8% relative improvement over the baseline’s 0.09. In the acoustic domain, our optimized Wav2Vec2.0 implementation reaches $\rho_{val}$=0.2734, outperforming the baseline (0.24) by 13.9%. This significant performance gain in audio processing suggests our feature enhancement strategy effectively captures subtle paralinguistic cues critical for emotional mimicry estimation.
5.2 Multimodal Fusion Effectiveness
Table 3 reveals the progressive performance enhancement through multimodal integration. Our key findings include:
  • The visual-textual pair achieves $\rho_{val}$=0.2934, surpassing audio-visual fusion (0.2846) by 3.1%
  • The audio-textual pair demonstrates strong synergy with $\rho_{val}$=0.3518, highlighting text modality’s complementary role
  • The full multimodal integration achieves state-of-the-art performance ($\rho_{val}$=0.3626), showing 27.5% improvement over best unimodal results
The progressive performance gains confirm our hypothesis about cross-modal complementarity. Notably, the textual modality contributes disproportionately to system performance, suggesting linguistic context plays a crucial role in disambiguating emotional mimicry patterns. Our dynamic fusion mechanism successfully preserves modality-specific information while enabling effective cross-modal interaction, as evidenced by the super-additive performance improvement in three-modality fusion.
Features Modality Mean $\rho_{val}$
ViT(baseline)[15] $V$ 0.09
ViT(ours) $V$ 0.16
Wav2Vec2.0(baseline)[15] $A$ 0.24
Wav2Vec2.0(ours) $A$ 0.2734
Table 2: The unimodal results on validation set of the EMI Estimation Challenge. We report the Pearson correlation coefficient ($\rho$) for the average of 6 emotion targets. Where $V$ represents the visual modality, and $A$ represents the acoustic modality.
Visual Audio Text Mean $\rho_{val}$
ViT Wav2Vec2.0 0.2846
ViT BERT 0.2934
Wav2Vec2.0 BERT 0.3518
ViT Wav2Vec2.0 BERT 0.3626
Table 3: The Multimodal results on validation set of the EMI Estimation Challenge.We use our pre-trained encoders for three modalities to conduct ablation experiments.
6 Conclusion
This paper presents our solution to the Emotional Mimicry Intensity (EMI) estimation challenge in the 8th Affective Behavior Analysis in-the-wild (ABAW) Competition. By introducing a dual-stage training paradigm and quality-aware dynamic modality fusion mechanism, our method achieves enhanced robustness in EMI estimation while maintaining tight three-modality alignment. Systematic validation on the Hume-Vidmimic2 dataset demonstrates that our approach attains an average Pearson correlation coefficient of 0.35 across six emotional dimensions, representing a 40% improvement over state-of-the-art baselines.